# Adaptive Regularization for Large-Scale Sparse Feature Embedding Models

### v1

@@ 4 Experimental Validation
To evaluate the effectiveness of our proposed method, we conduct several experiments on different public datasets and our industrial dataset. The experimental setup is described in section 4.1. The training framework proposed by zhu2022bars is used for training on public datasets, while XDL (jiang2019xdl) is employed for training our proprietary industrial dataset. We show the learning curve on the Avazu dataset in section 4.2 to demonstrate the better generalization of our methods over multi-epoch training. In section 4.3, we compare multiple backbones and datasets to prove the generalization of our method. Finally, we give a detailed example to show that the low-frequency embeddings are the primary contributors of one-epoch problem, which demonstrates we should allocate a smaller norm budget to embeddings with lower sample frequency. For reproduction, the code [^1] https://anonymous.4open.science/r/AdaptiveRegularization-AIDC/ is available.
@@ 4.1 Experimental Setup
4.1.1 Datasets
In our experiments, we use three public datasets as well as an online business dataset from sponsored search. Details of these datasets are provided below.
iPinYou [^2] https://huggingface.co/datasets/reczoo/iPinYou_x1. The training dataset includes processed bidding, impression, click, and conversion logs from iPinYou DSP. It contains 19.5 million records and 16 categorical features, providing a suitable context for exemplifying the overfitting phenomenon.
Amazon [^3] https://huggingface.co/datasets/reczoo/AmazonElectronics_x1. This is a widely used dataset from Amazon for evaluating CTR estimation models. In our study, we use the electronics category of the Amazon dataset, which contains approximately 3 million records and 3 categorical features. Popular models such as DIN (zhou2018deep) has used this dataset for performance evaluation.
Avazu [^4] https://huggingface.co/datasets/reczoo/Avazu_x1. This dataset comprises approximately 10 days of labeled click-through data from mobile advertisements, consisting of 40 million records and 22 categorical features spanning both user attributes and advertisement characteristics. Many of these features are highly dimensional, making this dataset a common benchmark for CTR estimation.
4.1.2 MLP Backbone
We test our method with four backbones as follows:
DNN. It is a basic deep learning model defined in equation 2.
WDL (cheng2016wide). It combines a linear model and a deep neural network to capture both memorization and generalization for recommender systems.
DeepFM (guo2017deepfm). It integrates factorization machines for learning low-order feature interactions and deep neural networks for high-order feature interactions.
WuKong (gu2022wukong). It proposes a specific model designed to explore these scaling laws. Its architecture is constructed with deeper and wider layers to capture high-order feature interactions.
4.1.3 Methods and Hyperparameters
To evaluate the one-epoch overfitting phenomenon, we set the number of training epochs to 4 and compare the performance of four benchmarks as follows:
Baseline optimizer. Adam and Adagrad serve as the baseline optimizers.
MEDA. MEDA is applied using the same optimizer settings as in baseline.
AdamW and AdagradW. The baseline optimizers are enhanced with weight decay.
AdamAR and AdagradAR. The baseline optimizers are combined with our method.
The hyperparameter settings are as follows. The embedding dimension is set to 32, with zero initialization. The learning rate is set to 0.001 for Adam and 0.01 for Adagrad, respectively. The batch size is 2048. Both $\alpha$ and the weight decay parameter are selected via grid search over the range $10^{-7}$ to $10^{-1}$, with the best hyperparameter settings determined according to validation performance (see appendix D for details). Network architectures are configured with minor adjustments across datasets. For detailed configurations, please refer to table 3 in appendix C. All other hyperparameters are set to their default values. Area under the curve (AUC) and binary cross-entropy loss are used as evaluation metrics. Experiments are conducted on a single machine with an NVIDIA L20 GPU.
@@ 4.3 Performance over Different Datasets and MLP Backbones
In this section, we compare test AUC across different datasets and MLP backbones to demonstrate the generalization capability of our method. The results, summarized in table 1 and table 2, show that our methods consistently outperform other approaches on all datasets and architectures, achieving superior AUC in both single-epoch and multi-epoch training. Furthermore, the gains are consistent across diverse model architectures, ranging from basic DNN to more sophisticated designs like WuKong, which capture complex feature interactions. Overall, these results demonstrate the robustness and versatility of our adaptive regularization approach across various settings.
Table 1: Comparison of test AUC across different datasets and models with the Adam optimizer. E1-E4 denote results after 1-4 epochs, respectively. The best results are highlighted in bold.
Dataset Method DNN WDL DeepFM WuKong
@@
iPinYou Adam 0.7498 0.7258 0.7174 0.7026 0.7622 0.7322 0.6972 0.6978 0.7595 0.7343 0.7057 0.6983 0.7644 0.7437 0.7137 0.6955
MEDA 0.7498 0.7631 0.7713 0.7692 0.7622 0.7582 0.7574 0.7506 0.7595 0.7582 0.7546 0.7518 0.7644 0.7671 0.7686 0.7714
AdamW 0.7486 0.7595 0.7628 0.7500 0.7609 0.7615 0.7692 0.7644 0.7600 0.7584 0.7486 0.7367 0.7593 0.7687 0.7586 0.7549
AdamAR 0.7549 0.7728 0.7730 0.7725 0.7675 0.7701 0.7692 0.7650 0.7677 0.7683 0.7637 0.7590 0.7668 0.7737 0.7722 0.7738
Amazon Adam 0.8482 0.8532 0.8314 0.8179 0.8489 0.8500 0.8255 0.8123 0.8496 0.8510 0.8236 0.8117 0.8574 0.8577 0.8355 0.8219
MEDA 0.8482 0.8537 0.8561 0.8574 0.8489 0.8506 0.8565 0.8557 0.8496 0.8511 0.8560 0.8557 0.8574 0.8577 0.8601 0.8627
AdamW 0.8472 0.8554 0.8409 0.8268 0.8473 0.8524 0.8377 0.8195 0.8472 0.8523 0.8380 0.8192 0.8544 0.8615 0.8442 0.8315
AdamAR 0.8506 0.8669 0.8709 0.8683 0.8513 0.8656 0.8690 0.8645 0.8513 0.8657 0.8690 0.8644 0.8576 0.8687 0.8702 0.8664
Avazu Adam 0.7483 0.7229 0.6973 0.6868 0.7485 0.7226 0.6949 0.6895 0.7488 0.7196 0.6918 0.6821 0.7479 0.7322 0.7000 0.7147
MEDA 0.7483 0.7508 0.7479 0.7479 0.7485 0.7475 0.7472 0.7516 0.7488 0.7478 0.7441 0.7523 0.7479 0.7538 0.7555 0.7559
AdamW 0.7584 0.7594 0.7579 0.7576 0.7586 0.7590 0.7544 0.7580 0.7590 0.7586 0.7546 0.7580 0.7511 0.7533 0.7565 0.7553
AdamAR 0.7621 0.7640 0.7636 0.7625 0.7630 0.7622 0.7617 0.7630 0.7629 0.7623 0.7613 0.7636 0.7619 0.7607 0.7624 0.7620
LZD Adam 0.7148 0.6661 0.6250 0.6105 0.7147 0.6719 0.6305 0.6053 0.7131 0.6706 0.6237 0.6105 0.7156 0.6629 0.6165 0.6069
MEDA 0.7148 0.7110 0.7113 0.7177 0.7147 0.7183 0.7169 0.7157 0.7131 0.7125 0.7157 0.7149 0.7156 0.7181 0.7182 0.7185
AdamW 0.7146 0.7132 0.7107 0.7138 0.7142 0.7136 0.7143 0.7129 0.7121 0.7107 0.7140 0.7132 0.7126 0.7149 0.7163 0.7115
AdamAR 0.7238 0.7256 0.7238 0.7234 0.7226 0.7233 0.7236 0.7242 0.7227 0.7199 0.7218 0.7221 0.7239 0.7218 0.7201 0.7185
Table 2: Comparison of test AUC across different datasets and models with the Adagrad optimizer. E1-E4 denote results after 1-4 epochs, respectively. The best results are highlighted in bold.
Dataset Method DNN WDL DeepFM WuKong
@@
iPinYou Adagrad 0.7626 0.6496 0.6366 0.6079 0.7623 0.6697 0.6424 0.6190 0.7626 0.6745 0.6440 0.6202 0.7651 0.6962 0.6489 0.5994
MEDA 0.7626 0.7673 0.7733 0.7689 0.7623 0.7725 0.7650 0.7717 0.7626 0.7722 0.7660 0.7728 0.7651 0.7756 0.7730 0.7712
AdagradW 0.7586 0.7614 0.7651 0.7570 0.7538 0.7652 0.7656 0.7598 0.7511 0.7662 0.7631 0.7632 0.7544 0.7704 0.7657 0.7645
AdagradAR 0.7717 0.7706 0.7759 0.7709 0.7705 0.7763 0.7758 0.7725 0.7703 0.7767 0.7741 0.7739 0.7689 0.7803 0.7792 0.7774
Amazon Adagrad 0.8435 0.8392 0.8127 0.8019 0.8427 0.8359 0.8113 0.7985 0.8425 0.8359 0.8102 0.8005 0.8487 0.8465 0.8176 0.8050
MEDA 0.8435 0.8477 0.8506 0.8521 0.8427 0.8480 0.8500 0.8503 0.8425 0.8493 0.8471 0.8522 0.8487 0.8568 0.8584 0.8613
AdagradW 0.8429 0.8564 0.8553 0.8533 0.8432 0.8554 0.8531 0.8492 0.8436 0.8556 0.8533 0.8492 0.8513 0.8616 0.8590 0.8584
AdagradAR 0.8474 0.8652 0.8709 0.8716 0.8462 0.8641 0.8696 0.8704 0.8462 0.8645 0.8697 0.8707 0.8539 0.8675 0.8708 0.8697
Avazu Adagrad 0.7547 0.7324 0.7154 0.7082 0.7555 0.7321 0.7175 0.7080 0.7553 0.7317 0.7165 0.7070 0.7531 0.7324 0.7170 0.7039
MEDA 0.7547 0.7544 0.7532 0.7548 0.7555 0.7547 0.7544 0.7556 0.7553 0.7545 0.7545 0.7550 0.7531 0.7543 0.7561 0.7547
AdagradW 0.7579 0.7586 0.7569 0.7563 0.7588 0.7572 0.7573 0.7593 0.7592 0.7586 0.7570 0.7586 0.7482 0.7554 0.7549 0.7551
AdagradAR 0.7626 0.7629 0.7629 0.7622 0.7629 0.7640 0.7632 0.7634 0.7628 0.7633 0.7626 0.7620 0.7621 0.7617 0.7620 0.7612
LZD Adagrad 0.7232 0.6666 0.6417 0.6400 0.7252 0.6696 0.6374 0.6388 0.7254 0.6677 0.6409 0.6399 0.7280 0.6753 0.6449 0.6365
MEDA 0.7232 0.7206 0.7137 0.7237 0.7252 0.7241 0.7253 0.7202 0.7254 0.7256 0.7245 0.7223 0.7280 0.7254 0.7252 0.7213
AdagradW 0.7190 0.7057 0.7071 0.7119 0.7210 0.7139 0.7142 0.7100 0.7209 0.7110 0.7098 0.7105 0.7179 0.7076 0.6834 0.6869
AdagradAR 0.7232 0.7261 0.7204 0.7265 0.7284 0.7261 0.7256 0.7269 0.7280 0.7267 0.7258 0.7263 0.7281 0.7258 0.7254 0.7213
@@ 4.4 Example of the Root Cause of Overfitting
In this section, we give an example to show the root cause of multi-epoch overfitting. We use the iPinYou dataset with DNN backbone and Adam to illustrate this issue. The detailed feature statistics are listed in table 4 in appendix E. We can observe that most features on iPinYou dataset are relatively dense, and their mean update intervals during training are less than 1. To investigate the root cause of overfitting, we select the feature ”IP”, which is the most sparse feature in iPinYou dataset with 704,966 unique IDs. We apply a filtering procedure to reduce its feature sparsity. Given a ratio $r$, we only retain the top-$r$ fraction (by frequency) of IDs and replace other IDs with a default ID.
Figure 2(a) demonstrates that as $r$ decreases, the test AUC remains stable across multiple epochs, indicating that the one-epoch overfitting phenomenon is effectively alleviated. It suggests that low-frequency IDs in the ”IP” feature are the primary cause of one-epoch overfitting in the iPinYou dataset. Based on this observation, a straightforward approach to mitigate this issue is to exclude the ”IP” feature. However, as shown by the case $r=0$, removing the ”IP” feature results in a substantial test AUC drop (from 0.7498 to 0.7429) at the end of the first epoch.

### v2

@@ 4 Experimental Validation
To evaluate the effectiveness of our proposed method, we conduct several experiments on different public datasets and our industrial dataset. The experimental setup is described in section 4.1. The training framework proposed by Zhu et al. (2022) is used for training on public datasets, while XDL (Jiang et al. (2019)) is employed for training our proprietary industrial dataset. We show the learning curve on the Avazu dataset in section 4.2 to demonstrate the better generalization of our methods over multi-epoch training. In section 4.3, we compare multiple backbones and datasets to prove the generalization of our method. We give a detailed example to show that the low-frequency embeddings are the primary contributors of the one-epoch problem in section 4.4, which demonstrates we should allocate a smaller norm budget to embeddings with lower sample frequency. Finally, in section 4.5, we conduct an ablation study to further clarify the contribution of occurrence interval estimation. For reproduction, the code [^2] https://anonymous.4open.science/r/AdaptiveRegularization-AIDC/ is available, and we use the same seed for identical experimental settings.
@@ 4.1 Experimental Setup
Datasets. In our experiments, we use three public datasets iPinYou [^3] https://huggingface.co/datasets/reczoo/iPinYou_x1, Amazon [^4] https://huggingface.co/datasets/reczoo/AmazonElectronics_x1, Avazu [^5] https://huggingface.co/datasets/reczoo/Avazu_x1 along with LZD, a proprietary online business dataset from sponsored search. Dataset details are provided in appendix L.
MLP Backbones. We evaluate our method using four MLP backbones, namely DNN, WDL (Cheng et al. (2016)), xDeepFM (Lian et al. (2018)) and WuKong (Zhang et al. (2024))
Methods. To evaluate the one-epoch overfitting phenomenon, we train the models for 4 epochs and compare 4 benchmarks. 1) Baseline optimizer: Adam and Adagrad serve as the baseline optimizers. 2) MEDA: MEDA is applied using the same optimizer settings as in baseline. 3) AdamW and AdagradW: The baseline optimizers are enhanced with weight decay only on embedding layers. 4) SAM: The baseline optimizers are combined with SAM (Foret et al. (2020)). 5) AdamAR and AdagradAR: The baseline optimizers are combined with our method.
Hyperparameters. The embedding dimension is set to 32, with zero initialization. The learning rate is set to 0.001 for Adam and 0.01 for Adagrad, respectively. The batch size is 2048. Both $\alpha$ and the weight decay parameter are selected via grid search over values of the form $10^{n}$ where $n$ ranges from 6.5-6.5 to 1-1 with step size of $0.5$. The optimal hyperparameters are selected based on validation performance (see appendix I for details). Network architectures are configured with minor adjustments across datasets. For detailed configurations, please refer to table 6 in appendix H. All other hyperparameters are set to their default values. Area under the curve (AUC) and binary cross-entropy loss are used as evaluation metrics. Experiments are conducted on a single machine with an NVIDIA L20 GPU.
@@ 4.3 Performance over Different Datasets and MLP Backbones
In this section, we compare test AUC across different datasets and MLP backbones to demonstrate the generalization capability of our method. Each experiment is repeated three times with different seed. We report the average test AUC using Adam and Adagrad optimizers in table 1 and table 2, while the standard deviation and detailed scalability comparison are presented in appendix J. The results show that our methods consistently outperform MEDA and weight decay method on all datasets and architectures in single-epoch training, except for the Amazon dataset which has less features and samples where SAM achieves superior performance. Notably, in multi-epoch training, our method achieves the highest AUC, surpassing all other approaches across every dataset and architecture. Furthermore, the gains are consistent across diverse model architectures, ranging from basic DNN to more sophisticated designs like WuKong, which capture complex feature interactions. Overall, these results demonstrate the robustness and versatility of our adaptive regularization approach across various settings.
Table 1: Comparison of average test AUC across different datasets and models using Adam optimizer. E1-E4 denote results after 1-4 epochs, respectively. The best results are highlighted in bold.
Dataset Method DNN WDL xDeepFM WuKong
@@
iPinYou Adam 0.7515 0.7304 0.7061 0.7014 0.7619 0.7320 0.7028 0.6987 0.7590 0.7391 0.6969 0.6844 0.7611 0.7442 0.7082 0.6915
MEDA 0.7515 0.7644 0.7684 0.7717 0.7619 0.7589 0.7565 0.7551 0.7590 0.7584 0.7575 0.7597 0.7611 0.7663 0.7662 0.7706
SAM 0.7510 0.7593 0.7445 0.7256 0.7610 0.7581 0.7404 0.7248 0.7565 0.7517 0.7413 0.7274 0.7033 0.7485 0.7465 0.7306
AdamW 0.7475 0.7592 0.7623 0.7568 0.7551 0.7656 0.7646 0.7634 0.7607 0.7660 0.7651 0.7574 0.7579 0.7634 0.7605 0.7511
AdamAR 0.7566 0.7692 0.7688 0.7724 0.7655 0.7729 0.7670 0.7668 0.7725 0.7733 0.7711 0.7678 0.7653 0.7736 0.7748 0.7736
Amazon Adam 0.8482 0.8548 0.8335 0.8180 0.8474 0.8510 0.8261 0.8156 0.8460 0.8535 0.8287 0.8163 0.8580 0.8600 0.8348 0.8232
MEDA 0.8482 0.8544 0.8556 0.8573 0.8474 0.8506 0.8566 0.8566 0.8460 0.8519 0.8551 0.8562 0.8580 0.8611 0.8621 0.8626
SAM $0.8507$ 0.8587 0.8417 0.8249 0.8516 0.8567 0.8396 0.8213 0.8505 0.8587 0.8390 0.8232 0.8588 0.8639 0.8404 0.8225
AdamW 0.8476 0.8571 0.8426 0.8276 0.8461 0.8533 0.8380 0.8223 0.8446 0.8557 0.8381 0.8240 0.8564 0.8632 0.8478 0.8349
AdamAR 0.8507 0.8683 0.8708 0.8686 0.8496 0.8654 0.8689 0.8659 0.8483 0.8676 0.8687 0.8675 0.8582 0.8696 0.8693 0.8664
Avazu Adam 0.7461 0.7205 0.7014 0.6883 0.7483 0.7221 0.6982 0.6886 0.7488 0.7217 0.7019 0.6869 0.7514 0.7360 0.7141 0.7079
MEDA 0.7461 0.7498 0.7489 0.7485 0.7483 0.7488 0.7480 0.7494 0.7488 0.7489 0.7505 0.7506 0.7514 0.7548 0.7553 0.7571
SAM 0.7451 0.7194 0.7013 0.6899 0.7477 0.7205 0.6993 0.6902 0.7484 0.7190 0.7010 0.6902 0.7513 0.7333 0.7155 0.7156
AdamW 0.7572 0.7582 0.7582 0.7583 0.7585 0.7581 0.7563 0.7570 0.7583 0.7582 0.7589 0.7593 0.7542 0.7547 0.7558 0.7564
AdamAR 0.7617 0.7631 0.7629 0.7629 0.7629 0.7629 0.7627 0.7626 0.7628 0.7633 0.7638 0.7636 0.7624 0.7612 0.7623 0.7624
LZD Adam 0.7118 0.6613 0.6252 0.6065 0.7155 0.6726 0.6308 0.6079 0.7164 0.6787 0.6321 0.6050 0.7101 0.6645 0.6102 0.6044
MEDA 0.7118 0.7105 0.7081 0.7176 0.7155 0.7162 0.7177 0.7162 0.7164 0.7170 0.7152 0.7139 0.7101 0.7170 0.7129 0.7128
SAM 0.7130 0.6696 0.6341 0.6155 0.7161 0.6795 0.6360 0.6111 0.7166 0.6750 0.6344 0.6134 0.7146 0.6713 0.6443 0.6212
AdamW 0.7132 0.7135 0.7140 0.7139 0.7143 0.7138 0.7142 0.7131 0.7142 0.7151 0.7139 0.7139 0.7115 0.7135 0.7152 0.7142
AdamAR 0.7229 0.7235 0.7241 0.7234 0.7233 0.7240 0.7246 0.7240 0.7244 0.7256 0.7242 0.7238 0.7227 0.7215 0.7208 0.7202
Table 2: Comparison of average test AUC across different datasets and models using Adagrad optimizer. E1-E4 denote results after 1-4 epochs, respectively. The best results are highlighted in bold.
Dataset Method DNN WDL xDeepFM WuKong
@@
iPinYou Adagrad 0.7593 0.6507 0.6231 0.6095 0.7646 0.6653 0.6321 0.6241 0.7674 0.6843 0.6505 0.6527 0.7661 0.6900 0.6497 0.6320
MEDA 0.7593 0.7686 0.7710 0.7729 0.7646 0.7681 0.7685 0.7715 0.7674 0.7722 0.7728 0.7740 0.7661 0.7705 0.7729 0.7695
SAM 0.7576 0.7661 0.7487 0.7303 0.7624 0.7676 0.7518 0.7369 0.7637 0.7661 0.7499 0.7363 0.7409 0.7678 0.7525 0.7427
AdagradW 0.7558 0.7651 0.7667 0.7595 0.7593 0.7675 0.7635 0.7593 0.7665 0.7673 0.7666 0.7652 0.7578 0.7661 0.7649 0.7600
AdagradAR 0.7681 0.7754 0.7760 0.7744 0.7731 0.7772 0.7748 0.7720 0.7760 0.7768 0.7762 0.7745 0.7718 0.7776 0.7782 0.7774
Amazon Adagrad 0.8438 0.8402 0.8141 0.8042 0.8406 0.8345 0.8085 0.7982 0.8405 0.8374 0.8103 0.7996 0.8491 0.8472 0.8156 0.8046
MEDA 0.8438 0.8481 0.8495 0.8505 0.8406 0.8470 0.8497 0.8510 0.8405 0.8463 0.8476 0.8500 0.8491 0.8569 0.8587 0.8609
SAM 0.8500 0.8527 0.8361 0.8193 0.8490 0.7995 0.8325 0.8155 0.8483 0.8521 0.8325 0.8170 0.8556 0.8567 0.8356 0.8174
AdagradW 0.8428 0.8578 0.8563 0.8535 0.8424 0.8553 0.8531 0.8498 0.8410 0.8565 0.8522 0.8508 0.8513 0.8630 0.8617 0.8606
AdagradAR 0.8479 0.8659 0.8711 0.8712 0.8453 0.8631 0.8687 0.8707 0.8444 0.8641 0.8681 0.8703 0.8538 0.8690 0.8708 0.8700
Avazu Adagrad 0.7541 0.7323 0.7164 0.7074 0.7543 0.7305 0.7158 0.7073 0.7550 0.7319 0.7160 0.7069 0.7548 0.7311 0.7160 0.7041
MEDA 0.7541 0.7549 0.7544 0.7545 0.7543 0.7547 0.7540 0.7551 0.7550 0.7538 0.7550 0.7553 0.7548 0.7551 0.7556 0.7564
SAM 0.7553 0.7333 0.7199 0.7094 0.7557 0.7328 0.7178 0.7079 0.7564 0.7332 0.7198 0.7094 0.7558 0.7353 0.7211 0.7110
AdagradW 0.7578 0.7580 0.7575 0.7566 0.7584 0.7581 0.7567 0.7585 0.7583 0.7581 0.7580 0.7585 0.7524 0.7555 0.7555 0.7563
AdagradAR 0.7628 0.7629 0.7630 0.7624 0.7631 0.7634 0.7623 0.7626 0.7636 0.7633 0.7635 0.7633 0.7628 0.7626 0.7627 0.7620
LZD Adagrad 0.7226 0.6658 0.6433 0.6376 0.7244 0.6695 0.6389 0.6386 0.7222 0.6675 0.6400 0.6382 0.7247 0.6726 0.6412 0.6339
MEDA 0.7226 0.7183 0.7150 0.7236 0.7244 0.7241 0.7239 0.7219 0.7222 0.7237 0.7253 0.7256 0.7247 0.7237 0.7234 $0.7239$
SAM 0.7213 0.6761 0.6446 0.6222 0.7229 0.6839 0.6481 0.6266 0.7229 0.6854 0.6480 0.6268 0.7248 0.6802 0.6479 0.6299
AdagradW 0.7145 0.7132 0.7135 0.7154 0.7138 0.7149 0.7137 0.7145 0.7155 0.7135 0.7117 0.7150 0.7127 0.7111 0.7147 0.7136
AdagradAR 0.7253 0.7247 0.7234 0.7241 0.7269 0.7258 0.7251 0.7257 0.7273 0.7268 0.7264 0.7263 0.7269 0.7260 0.7252 0.7239
@@ 4.4 Example of the Root Cause of Overfitting
In this section, we give an example to show the root cause of multi-epoch overfitting. We use the iPinYou dataset with DNN backbone and Adam to illustrate this issue by single experiment. The detailed feature statistics are listed in table 7 in appendix K. We can observe that most features on iPinYou dataset are relatively dense. To investigate the root cause of overfitting, we select the feature ”IP”, which is the most sparse feature in iPinYou dataset with 704,966 unique IDs. We apply a filtering procedure to reduce its feature sparsity. Given a ratio $r$, we only retain the top-$r$ fraction (by frequency) of IDs and replace other IDs with a default ID.
Figure 2(a) demonstrates that as $r$ decreases, the test AUC remains stable across multiple epochs, indicating that the one-epoch overfitting phenomenon is effectively alleviated. It suggests that low-frequency IDs in the ”IP” feature are the primary cause of one-epoch overfitting in the iPinYou dataset. However, as shown by the case $r=0$, removing the ”IP” feature results in a substantial test AUC drop (from 0.7498 to 0.7429) at the end of the first epoch.
@@
4.5 Ablation Study and Bucket Analysis
We use the iPinYou dataset for analysis because it contains a single feature causing the one-epoch issue as described in section 4.4. Since ”IP” can be interpreted as a proxy for a user, we create 5 buckets based on the frequency of the ”IP” feature to perform user-based bucket analysis, examining AUC gains, regularization strength, and $\ell_{2}$ norm of our method. A smaller bucket index indicates a lower occurrence frequency. Figure 3 shows that the bucket norms can be controlled via adaptive regularization strength while preserving the AUC gains across all buckets, and our method achieves particularly strong performance in the high-frequency bucket due to the larger norm budget available. Then, we conduct an ablation study on occurrence interval estimation with a DNN backbone. In the AdamW baseline, we apply a constant weight decay to the embedding layers only. We then gradually apply our method to different buckets from 1 to 5. Table 3 shows that, compared with using a constant weight decay, decreasing the weight decay for high-frequency features and increasing it for low-frequency features can further improve performance while alleviating the one-epoch issue.
(a) Test AUC
(b) Average regularization strength
(c) $\ell_{2}$ norm
Figure 3: Bucket analysis of ”IP” feature on the iPinYou dataset with DNN backbone and Adam optimizer. (a) shows the test AUC results over different feature frequency buckets at the end of epoch 2. (b) presents the regularization strength. (c) shows the cumulative $\ell_{2}$ norm of ”IP” feature at the end of epoch 2.
Table 3: Comparison of test AUC for ablation study using Adam optimizer.
Experiment Setting E1 E2 E3 E4
AdamW 0.7486 0.7595 0.7628 0.7500
AdamAR-Bucket 1 & AdamW-Bucket 2-5 0.7457 0.7607 0.7646 0.7520
AdamAR-Bucket 1-2 & AdamW-Bucket 3-5 0.7496 0.7622 0.7655 0.7556
AdamAR-Bucket 1-3 & AdamW-Bucket 4-5 0.7510 0.7646 0.7657 0.7614
AdamAR-Bucket 1-4 & AdamW-Bucket 5 0.7470 0.7656 0.7660 0.7635
AdamAR 0.7549 0.7728 0.7730 0.7725